A company wants to raise $500 million in a new stock issue. Its investment banker indicates that the sale of new stock will require 8 percent underpricing and a 7 percent spread.   It’s current stock price is $75. Round answers & intermediate calculations to the ones place. Spread is calculated on the amount after the underpricing.

What will be the price to the public? $

What is the net amount per share received by the company? $

How many shares must the company sell?

How much spread will the investment banker earn on the sale? $

Does the company lose any cash flow during the issuance due to underpricing? Enter yes or no
Answer #1

Price to public = Current price - underpricing

= 75 - 8%*75 = $69

Net amount received by Company = 69 - 7%*69

= $64.17

Shares required to be sold = Amount required/Amount received per share

= 500,000,000/64.17

= 7,791,803.02 shares

i.e. 7,791,803 shares

Spread earned = 7,791,803*4.83 = $37,634,408.49

Yes